OpenStudy (anonymous):

A building measures 50 feet in length and w feet in width. Which of the following is an expression for the perimeter of the building? A. 50/w B. 50w C. 50 + w D. 100 + 2w

OpenStudy (anonymous):

P = 2l + 2w P = 2(50) + 2w P = 100 +2w Answer D.